OpenStudy (anonymous):

The sides of a polygon are 3, 5, 6, 8, and 10. The perimeter of a similar polygon is 40. Find the sides of the second polygon (from smallest to largest). Put answers in decimal form (hundredths).

OpenStudy (anonymous):

First find the factor by which every side is multiplied to get the similar polygon, you can do this by letting the factor = x and solve... 3x + 5x + 6x + 8x + 10x = 40 32x = 40 x = 5/4 =1/25 now just multiply every number by 1.25 to get the answer

OpenStudy (anonymous):

3.75,6.25,7.5,10,12.5

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

Given: rectangle ABCD is similar to rectangle ZBXY. Enter answer as a simplified, improper fraction. If ZY = 5, XC = 3, DC = 4, then XY =